OpenStudy (anonymous):

A hot Wheel is rolled off a table 85 cm tall and lands 75 cm from the edge or the table. How fast was the Hot Wheel racing?

OpenStudy (anonymous):

Hmm \[t = \sqrt{\frac{2h}{g}}\] \[t = \sqrt{.17}\] \[x = v\times t\] \[v =\frac{x}{t}\] \[v = \frac{.75}{\sqrt{.17}}\] I love Hotwheels

OpenStudy (anonymous):

thanks bro